OpenStudy (anonymous):

How can x + 3 = -2x + 5 be set up as a system of equations?

OpenStudy (anonymous):

y = x + 13 y = -2x + 5 y = x + 3 y = -3x + 11 y = x + 12 y = -2x + 1 y = x + 3 y = -2x + 5